OpenStudy (anonymous):

A tile factory earns money by a flat fee for delivery and a sales price per tile sold. For one customer, the tile factory shipped 10,000 units to a retail store at a rate of $0.25 per tile. The total the tile factory earned from that sale was $3,000. Which of the following equations describes the revenue of the tile factory for this sale in terms of tiles sold?

OpenStudy (matt101):

Our equation will be in this form: R = pt + d R is the total revenue (money made) p is the price per tile t is the number of tiles sold d is the delivery charge This is a y = mx + b equation, I just switched the variables to make it easier to see where they come from. In y = mx + b equations, you know that you have numbers instead of m and b; it's only x and y that remain as variables. In our case, we only want t and R to remain variables (which makes sense, because the money made by the company depends on how many tiles were sold - the price of tiles and the delivery fee are always the same). Even though we know p from the question, to make this work, we still need to find d. We can use the numbers given in the equation to do this. R = pt + d 3000 = 0.25(10000) + d d = 500 Therefore, our general equation for revenue will be: R = 0.25t + 500